### v1

# Ego-Vision World Model for Humanoid Contact Planning

Abstract
Enabling humanoid robots to exploit physical contact, rather than simply avoid collisions, is crucial for autonomy in unstructured environments. Traditional optimization-based planners struggle with contact complexity, while on-policy reinforcement learning (RL) is sample-inefficient and has limited multi-task ability. We propose a framework combining a learned world model with sampling-based Model Predictive Control (MPC), trained on a demonstration-free offline dataset to predict future outcomes in a compressed latent space. To address sparse contact rewards and sensor noise, the MPC uses a learned surrogate value function for dense, robust planning. Our single, scalable model supports contact-aware tasks, including wall support after perturbation, blocking incoming objects, and traversing height-limited arches, with improved data efficiency and multi-task capability over on-policy RL. Deployed on a physical humanoid, our system achieves robust, real-time contact planning from proprioception and ego-centric depth images. Website: https://ego-vcp.github.io/
I Introduction
Humanoids are expected to advance from dynamic locomotion [1, 2] to intelligent interaction in complex, unstructured environments. Achieving this requires purposeful contact exploitation rather than simple avoidance. Effective humanoids must use their bodies to interact with the world as humans do, such as bracing against a wall for balance, blocking objects for safety, or ducking under obstacles. These contact-aware skills are essential for greater autonomy, robustness, and physical intelligence in robots.
Reasoning about contact remains challenging for humanoids [3, 4, 5]. Traditional optimization-based methods [6, 7] struggle with the combinatorial complexity of real-time contact scheduling and are sensitive to model inaccuracies, reducing adaptability to unforeseen situations. Parallelized simulation [8] has enabled on-policy RL to succeed in robot control for quadrupeds [9], bipeds [10], and humanoids [11]. However, these methods are sample-inefficient, especially with visual inputs [9], and struggle with multi-task learning and complex scene interactions.
Figure 1: An illustration of our framework in the “Support the Wall” task. When subjected to a sudden perturbation (left), the robot uses its learned world model to predict and plan a stabilizing action within its planning horizon (center). This allows it to successfully execute the plan and brace its hands against the wall to make contact and maintain balance (right).
We address this by integrating a learned world model with sampling-based Model Predictive Control (MPC). Our approach trains a scalable world model from a random, demonstration-free offline dataset, predicting future outcomes in compressed latent space rather than raw pixels, and understanding action consequences. We introduce a surrogate value function to guide planning, allowing MPC to efficiently evaluate candidate action sequences. This synergy enables agile, vision-based contact planning for humanoids across tasks with improved data efficiency and performance.
The main contributions of this work are as follows:
  1. 1.
    A Scalable Visual World Model for Dynamic Robots: We learn a visual world model that scalably captures the dynamics of diverse contact tasks, trained entirely on a demonstration-free offline dataset.
  2. 2.
    Planning from Pixels with Value-Guidance: We introduce a sampling-based MPC framework using a learned surrogate value function to guide the planning process.
  3. 3.
    Agile and Robust Real-world Visual Contact Planning: We validate the proposed framework on a physical humanoid robot, demonstrating multiple novel agile and robust contact planning tasks solely from ego-centric depth images and proprioceptive feedback.
Figure 2: World Model Training Pipeline. The pipeline begins with the offline data collection process shown in (a), where a dataset $\mathcal{D}$ of trajectories is generated by applying randomly sampled high-level actions (end-effector position $p_{ee}^{\top}$ and body height $h_{body}$) to a simulated humanoid equipped with a trained low-level policy. This dataset is then used to train the world model, as depicted in (b). At each timestep $t$, an observation $o_{t}$, consisting of a depth image and proprioception, is encoded into a stochastic latent state $z_{t}$, which is then decoded to produce a reconstruction $\hat{o}_{t}$. Concurrently, a recurrent network updates its latent state $h_{t}$ based on the previous state and action. From the combined latent state $(h_{t},z_{t})$, the model predicts (i) $\hat{z}_{t}$, a prior sample of the stochastic latent state; (ii) $\hat{d}_{t}$, the termination probability; and (iii) $\hat{Q}_{t}$ a surrogate action-value guiding the planner in evaluating different actions. All of these predictions are optimized against the ground-truth data from the offline trajectories, enabling the model to learn both the environment’s dynamics and a robust value function for planning.
II Related Work
II-A Model-Based Contact Planning
For both locomotion and manipulation, robotics is replete with contact-rich problems, made challenging by the non-smooth dynamics induced by the impact  [1]. Optimization-based approaches address this by explicitly modeling these physical interactions, like linearizing the complex friction model into a Linear Complementarity Problem (LCP) [12], or relaxing it into a Cone Complementarity Problem (CCP) [13]. These formulations can then be embedded within a trajectory optimization framework [6, 7]. Another prominent paradigm is Hybrid Zero Dynamics (HZD) [14, 15], which addresses the non-smooth contact dynamics of legged locomotion by enforcing virtual constraints whose associated zero dynamics surface remains invariant through impacts. However, such model-based approaches are often hindered by model inaccuracies and high computational costs [16], which complicate real-time deployment. Furthermore, their reliance on predefined structures, such as periodic gaits [17] or reference foot-end trajectories [14], makes it difficult to scale them to more general, aperiodic whole-body contact scenarios.
II-B Learning-Based Contact Planning
Learning-based approaches for real-world contact planning have shown remarkable potential, enabling dynamic skills [10, 9, 3, 5, 18, 19]. However, three significant challenges remain largely unaddressed. First, interaction with both dynamic and static objects is limited. Most existing work on legged locomotion based on simplified 2.5D elevation maps [4], which cannot represent dynamic or overhanging obstacles such as a moving ball or an archway. Second, sample efficiency remains a bottleneck. Modern approaches rely heavily on synthetic data [8], but the computational cost of rendering visual input makes on-policy RL algorithms prohibitively expensive [9]. The sparse and discontinuous nature of contact events also poses significant challenges for model-free methods, which often require extensive training or carefully designed inductive biases to guide exploration effectively [5, 20]. Finally, multi-task capability is limited. While current policies can be trained to solve a single task with a specific reward function, they often fail to generalize across interactions with different objects or adapt to variations in task definitions [21].
II-C Planning with Robotic World Models
The world model [22] is a learned, internal model of an environment that allows an robot to predict future outcomes based on actions. The emphasis is not on explicitly predicting the future, but rather on predicting its abstract representation. Using a learned world model for model-based planning offers a path toward better generalization and data efficiency, presenting a promising solution for contact planning [23, 24].
Early work on world models [22] was defined as enabling efficient policy learning in reinforcement learning [23, 24]. Today, the definition has evolved as a generative AI system [25, 26] capable of understanding and simulating the physical world’s causal relationships, dynamics, and interactions. In control and robotics, a similar path is being explored by learning neural dynamics models to represent complex systems [27, 28, 29, 30]. These learned models could be integrated into frameworks such as sampling-based MPC [31, 32, 16] to achieve highly adaptive control [33, 4, 34].
Nonetheless, enabling robotic world models to fully generalize remains an open problem. This is especially the case for contact planning, as the underlying whole-body contact state is not directly observable and is difficult to infer and predict from partial, noisy sensory data.
III Methods
This section begins by outlining our hierarchical control framework in Section III-A, composed of a low-level policy and our high-level planner. The subsequent sections then detail the two components of this high-level planner: the world model’s architecture and training (Section III-B), and the value-guided sampling MPC (Section III-C).
Figure 3: Value-Guided Sampling MPC. This figure illustrates how the trained world model is used for planning via value-guided sampling MPC. This process performs open-loop prediction to find the best action sequence starting from a single real observation. At inference time, this process begins by encoding the current observation $o_{t}$ into its latent state $z_{t}$, after which the planner samples a batch of $M=1024$ candidate action sequences over a planning horizon of $N=4$ steps. The world model predicts the future latent state ($h_{t+k},\hat{z}_{t+k}$) by recursively applying its learned dynamics model. At each prediction step, the surrogate value ($\hat{Q}_{t+k}$) evaluates the sampled actions, while the termination signal, $\hat{d}_{t}$, predicts the probability of robot failure, such as falling; if this probability exceeds a threshold of 0.9, all subsequent value estimates, $\hat{Q}$, for that trajectory are set to zero. The planner evaluates $M$ candidate trajectories, where the score for each trajectory is calculated by the objective function $\hat{J}_{N}$ in Eq. (16). This set of scored trajectories is then optimized using the Cross-Entropy Method (CEM) to find the optimal action sequence.
III-A Hierarchical Framework and Data Collection
Our framework consists of two components: a low-level whole-body policy capable of tracking diverse commands, and a high-level planner whose action specifies which commands to track, as shown in Fig. 2.
The low-level controller tracks commands $c=\left[v^{\top},p_{ee}^{\top},h_{body}\right]^{\top}$, where $v$ is the desired linear velocity, $p_{ee}$ the desired end-effector position, and $h_{body}$ the body height, while maintaining balance. Its observation space is purely proprioceptive, including angular velocity ($\omega$), the projected gravity vector ($g$), the command ($c$), joint positions ($q$), and joint velocities ($\dot{q}$). The controller is represented by a deep neural network and is trained in simulation using PPO, following established approaches [8, 35].
Once a reliable low-level controller is obtained, the high-level planner is designed to account for both contact and task planning. It specifies actions $a_{t}=\left[p_{ee}^{\top},h_{body}\right]^{\top}$ to the low-level controller based on an observation space $o_{t}$ combining proprioception (as in the low-level controller) with visual input from a downsampled 64×48 ego-centric depth image. We exclude base velocity $v$ from the planner’s action space to force the robot to solve contact-rich problems through postural manipulation, rather than learning locomotion-based evasion strategies. The high-level planner consists of a world model and a sampling-based MPC framework. We train the world model from the offline dataset and, at inference time, use it to evaluate $M$ candidate action sequences over an $N$-step horizon and optimize the action $a_{t}$.
Our offline dataset, $\mathcal{D}$,  is generated by collecting trajectories, $\tau$, from the simulated humanoid. At each timestep, the robot receives an observation $o_{t}$, executes a randomly sampled action $a_{t}$, and in return, receives the reward $r_{t}$ and termination signal $d_{t}$ from the environment. These collected transition tuples $\{o_{t},a_{t},r_{t},d_{t}\}$ are then stored in a final trajectory dataset structured as [Batch, Time, Data]. At each step, we sample the finite differences of planner actions $a_{t}=a_{t-1}+\eta\cdot\delta$ from $\delta\sim\mathcal{U}(-1,1)$, where $\eta$ is a scalar step that controls the magnitude of the update. This step is performed after normalizing the task space $\left[p_{ee}^{\top},h_{body}\right]^{\top}$ of the low-level controller. $\eta$ is set to $0.32$. The purpose of using such a method to sample the actions is to (1) avoid using any demonstration, which is expensive to obtain for the whole-body commands of a humanoid, and (2) heuristically avoid ineffective and jitter behavior data far outside the command range.
Figure 4: Real-World experiments validating the proposed framework. (a) A demonstration of sequential task execution and generalization, where the robot traverses an arch (i) and then blocks a previously unseen box (ii). (b) Support the wall to maintain balance by bracing the wall with the hands when pushed towards the wall. (c) Blocking both an in-distribution ball (with a size consistent with the training data) and an unseen box; (d) Squat and traverse an arch.
III-B Ego-Vision Humanoid World Model
Prior auto-regressive models learn system dynamics by mimicking existing controllers for continuous tasks like velocity tracking [29]. However, when applied to high-dimensional image observations, this pixel-prediction approach suffers from compounding errors over long horizons. This issue is exacerbated in contact-aware scenarios where defining a goal trajectory in pixel space is often intractable.
To address this, we draw inspiration from general world models like Dreamer [36] and JEPA [25]. We focus on predicting abstract latent of future observations, enabling the model to capture more fundamental structures within the data. As illustrated in Fig. 2, our world model is composed of several key components detailed below.
First, the world model leverages a recurrent neural network (RNN) to maintain a deterministic dynamics latent state $h_{t}$, which summarizes dynamics information across time. At each step, a stochastic latent state $z_{t}$, which extracts the abstract latent of the current observation, is inferred from the current observation $o_{t}$ and the latent state $h_{t}$. Similar to an autoencoder, the model is trained to reconstruct the observation $o_{t}$ as $\hat{o}_{t}$ after passing it through a latent bottleneck, which compels the latent state $z_{t}$ to encode the most salient and abstract features. For notational simplicity, we let $\phi$ denote the parameters of all world model components, with $q_{\phi}$ and $p_{\phi}$ representing the encoder and decoder, respectively. The overall process can then be expressed as:
$h_{t}$ $:=f_{\phi}(h_{t-1},z_{t-1},a_{t-1})$ (1)
$z_{t}$ $\sim q_{\phi}(z_{t}\mid h_{t},o_{t})$ (2)
$\hat{o}_{t}$ $\sim p_{\phi}(\hat{o}_{t}\mid h_{t},z_{t}).$ (3)
We assume $z_{t}$ and $\hat{o}_{t}$ to follow the Gaussian distribution.
In addition, we introduce a model that estimates the stochastic latent without using the current observation: given $h_{t}$, it yields a latent $\hat{z}_{t}\sim p_{\phi}(\hat{z_{t}}\mid h_{t})$ that closely approximates $z_{t}$, thereby enabling rollouts in latent space.
Different from Dreamer [36], we need to consider an architecture that better addresses the challenges unique to robotics in the real world, such as (1) significant partial observability, (2) high sensor noise, and (3) sparse contact. These factors make it difficult to predict contact-aware rewards from the observation. Therefore, we design specialized heads that, conditioned on the latent state $(h_{t},z_{t})$ and a candidate action $a_{t}$, directly estimate the expected long-term outcome. Specifically, we predict a termination probability $\hat{d}_{t}$ and a surrogate value $\hat{Q}_{t}$, which represents the expected cumulative return. This allows the robot to evaluate the potential response to different actions directly from its learned latent state.
$\hat{d}_{t}$ $\sim D_{\phi}(\hat{d}_{t}\mid h_{t},z_{t}),$ (4)
$\hat{Q}_{t}$ $:=Q_{\phi}(h_{t},z_{t},a_{t}).$ (5)
Our surrogate value could condition on the latent state $z_{t}$, allowing the robot to infer the current task context from its observations and dynamically adapt its objective. This enables us to train the model directly on a mixed dataset containing data from all tasks. We opt for a computationally efficient design consisting of a CNN for image feature extraction and MLPs for all other components.
The model is optimized by minimizing the single-step loss as shown below. This total loss is a simple sum of three main components: a reconstruction loss ($\mathcal{L}_{\text{rec}}$), a joint-embedding predictive loss ($\mathcal{L}_{\text{jep}}$), and a Q-loss ($\mathcal{L}_{\hat{Q}}$):
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{jep}}+\mathcal{L}_{\hat{Q}}$. (6)
The reconstruction loss, $\mathcal{L}_{\text{rec}}$, ensures the world model can extract a tight latent space of the environment. It is defined as a combination of a Negative Log-Likelihood (NLL) for the observation reconstruction ($\mathcal{L}_{\text{obs}}$) and a Binary Cross-Entropy (BCE) loss for the termination signal ($\mathcal{L}_{\text{term}}$):
$\mathcal{L}_{\text{rec}}=\mathbb{E}_{\hat{\mathbf{z}}_{t}\sim q_{\phi}(\mathbf{z}_{t}|\mathbf{o}_{t},\mathbf{h}_{t})}\left[\mathcal{L}_{\text{obs}}+\mathcal{L}_{\text{term}}\right].$ (7)
The joint-embedding predictive loss, $\mathcal{L}_{\text{jep}}$, consists of two KL divergence terms that enforce a consistent and non-collapsing latent space [37, 36].
$\mathcal{L}_{\text{jep}}=D_{\text{KL}}\Big(\text{sg}\big(q_{\phi}(\mathbf{z}_{t}|\mathbf{o}_{t},\mathbf{h}_{t})\big)\ \big|\big|\ p_{\phi}(\mathbf{z}_{t}|\mathbf{h}_{t})\Big)$
$+D_{\text{KL}}\Big(q_{\phi}(\mathbf{z}_{t}|\mathbf{o}_{t},\mathbf{h}_{t})\ \big|\big|\ \text{sg}\big(p_{\phi}(\mathbf{z}_{t}|\mathbf{h}_{t})\big)\Big),$ (8)
where sg is the stop-gradient operator.
The surrogate value loss, $\mathcal{L}_{\hat{Q}}$, is a mean-squared error term that trains the value function to estimate the target $Q_{\text{target}}$. We simply apply a Monte Carlo (MC) estimator here to get $Q_{\text{target}}$, and we empirically found that using an MC estimator yields more stable results in our scenario than TD-error:
$\mathcal{L}_{\hat{Q}}=\mathbb{E}_{\mathbf{\tau}\sim\mathbf{D}}\sum_{t}\mathbb{E}_{\mathbf{z}_{t}\sim q_{\phi}(\cdot|\mathbf{o}_{t},\mathbf{h}_{t})}\left[\left({Q}_{\phi}(\mathbf{h}_{t},\mathbf{z}_{t},\mathbf{a}_{t})-Q_{\text{target}}\right)^{2}\right].$ (9)
Figure 5: Sample efficiency comparison: Our method uses an offline dataset collected from random action, while PPO collects data from environments at every iteration. The x-axis represents the number of step transitions used, while the y-axis shows the reward for each task. A greater value on the x-axis indicates a larger amount of data used, and a higher value on the y-axis signifies better performance. While our method utilizes a dataset of at most 1M steps, we continued to train PPO for a greater number of steps to determine when it could achieve comparable performance.
III-C Value-Guided Sampling MPC
In practical applications involving complex robotics and perception, learning a perfect, optimal state-action value function and greedily maximizing it remains a significant challenge, which is further compounded when extending to observation-action value functions. This difficulty stems from two primary sources:
  1. 1.
    Challenges in Offline Learning: Finite offline datasets provide incomplete coverage, leading to unreliable value estimates for out-of-distribution actions.
  2. 2.
    Partial Observability and Physical Non-idealities: Robotic systems suffer from partial observability, as the full state, such as contact forces, is not observable. is not directly measured, and is subject to sensor noise and action delays, both of which degrade value estimation.
To address this, we introduce a Value-Guided Sampling MPC framework. This approach explicitly treats the learned value function not as an optimal oracle, but as a powerful, albeit imperfect, heuristic to guide a robust, receding-horizon planning process.
Consider a standard MDP with state $s_{t}$ and action $a_{t}$ for the analysis of the variance reduction of our proposed method. We denote the estimation of the true $Q$ functions as $\hat{Q}$. When the $Q$ is available, we can leverage the Bellman principle to obtain the optimal control input:
$\pi(s_{t})=\arg\max_{a_{t}}{Q}(s_{t},a_{t}).$ (10)
However, in practice, only the estimation $\hat{Q}$ is available, which may have a large variance. The imperfection of $\hat{Q}$ may lead to performance degradation. To mitigate this issue, instead of directly optimizing on the estimated Q function, we consider an $N$-step surrogate optimization:
$\hat{\pi}(s_{t})=\arg\max_{\{a_{0:N-1}\}}$ $\hat{J}_{N}=\frac{1}{N}\sum_{t=0}^{N-1}\hat{Q}(s_{t},a_{t})$ (11)
$\mathrm{s.t.}$ $\quad s_{t+1}=f(s_{t},a_{t})$
[…]
Consider the residual between $\hat{Q}$ and $Q$ as $\epsilon_{Q}$, and
$\hat{Q}(s_{t},a_{t})=Q(s_{t},a_{t})+\epsilon_{Q}(s_{t},a_{t}).$
For shorthand notation, we use the subscript $(t)$ to denote the quantity at time step $t$. We have the surrogate loss function:
$\frac{1}{N}\sum_{t=0}^{N-1}\hat{Q}_{t}=\frac{1}{N}\sum^{N-1}_{t=0}\left(Q_{t}+\epsilon_{t}\right).$ (12)
As we use Monte Carlo estimation to compute $\hat{Q}$ as an unbiased estimation, the mean of $\epsilon_{t}$ is zero, and thus the variance on the RHS can be obtained by
$\operatorname{Cov}[\hat{J}_{N}]=\operatorname{Var}[\frac{1}{N}\sum^{N-1}_{t=0}\epsilon_{t}]=\frac{1}{N^{2}}\sum_{(i,j)}\operatorname{Cov}[\epsilon_{i},\epsilon_{j}].$ (13)
We assume that the variance of $\epsilon_{t},\forall t,$ satisfies:
$0\leq V_{\min}\leq\operatorname{Var}[\epsilon_{t}]\leq V_{\max}$ (Bounded Variance)
and the correlation is bounded by $\rho<1$, such that $\forall i,j$
$|\operatorname{Cov}[\epsilon_{i},\epsilon_{j}]|\leq\rho\operatorname{Var}[\epsilon_{i}]\operatorname{Var}[\epsilon_{j}].$ (Bounded Correlation)
The upper bound variance is achieved if all steps are positively correlated:
$\operatorname{Cov}[\hat{J}_{N}]\leq\frac{N+\rho(N^{2}-N)}{N^{2}}V_{\max}=:V_{\operatorname{ub}}.$ (14)
For the lower bound, similarly, we have:
$\operatorname{Cov}[\hat{J}_{N}]\geq\max\left\{\frac{NV_{\min}-\rho(N^{2}-N)V_{\max}}{N^{2}},0\right\}=:V_{\operatorname{lb}}.$ (15)
Thus, we have limit when $\rho\rightarrow 0:V_{\operatorname{ub}}\rightarrow\frac{V_{\max}}{N},V_{\operatorname{lb}}\rightarrow\frac{V_{\min}}{N}.$ As we compute $\hat{Q}$ from an offline dataset generated by random action in each step, the correlation between $\hat{Q}_{i}$ from different time steps is weaker than by sampling using a state feedback policy. Thus, it is possible that the $\rho$ is small and thus $\hat{J}_{N}$ has a substantially lower variance than $\hat{Q}$. Although additional bias is introduced because the surrogate loss does not preserve the local optimum, if the variance of $\hat{Q}$ dominates, this strategy can significantly improve the performance.
Based on the above analysis, we apply the surrogate objective function $\hat{J}_{N}$ from  (11) using the latent variables $h_{t}$ and $\hat{z}$ as the representation of the robot states. The optimal sequence, denoted $A_{t}^{*}=\{a_{t}^{*},a_{t+1}^{*},...,a_{t+N-1}^{*}\}$, is the one that maximizes our surrogate objective $\hat{J}_{N}$:
$A_{t}^{*}=\arg\max_{A_{t}}$ $\ \ \frac{1}{N}\sum_{k=0}^{N-1}Q_{\phi}(h_{t+k},\hat{z}_{t+k},a_{t+k}),$ (16)
$\mathrm{s.t.}$ $\quad h_{t+k+1}=f_{\phi}(h_{t+k},\hat{z}_{t+k},a_{t+k}),$
$\quad\hat{z}_{t+k}\sim p_{\phi}(\hat{z}_{t+k}\mid h_{t+k}).$
As shown in Fig. 3, the framework’s memory is maintained in two latent states: a dynamics latent state $h_{t}$ and a current observation latent state $z_{t}$. The $h_{t}$ is computed by the RNN from the previous latent ($h_{t-1},z_{t-1}$) and the last action ($a_{t-1}$). The $z_{t}$ is then generated in one of two ways: when an observation $o_{t}$ is available, $z_{t}$ is inferred using both the latent state $h_{t}$ and the observation $o_{t}$; for future predictions, it is generated from the latent state $h_{t}$ alone.
Our world model also predicts the probability of robot failure, $\hat{d}_{t}$, such as falling; if this probability exceeds a threshold of 0.9, all subsequent value estimates for that trajectory are set to zero.
We use the Cross-Entropy Method (CEM) to find optimal action sequence $A_{t}^{*}$. Once identified, only the first action $a_{t}^{*}$ is executed. This iterative re-planning allows the robot to continuously incorporate feedback from the environment, enabling it to react to disturbances and correct for model inaccuracies in real-time. From Table I, we select a planning horizon $N=4$ as our default, as it provides the best overall performance across all tasks.
Figure 6: Multi-task performance and latent space visualization. (a) A single model trained jointly on all tasks achieves comparable normalized performance to specialized single-task models. (b-c) The t-SNE visualizations reveal a clear separation of tasks in the latent spaces. As envisioned by our design, the latent $h_{t}$ primarily represents dynamics, showing significant evolution over time, while the latent $z_{t}$ provides a more compressed representation of the current environmental observation.
IV Experiments
Our experimental platform is the Unitree G1 humanoid robot, equipped with a RealSense D435i camera. All quantitative analyses, including ablation studies and baseline comparisons, are conducted in a controlled simulation environment. All comparisons are conducted using a consistent set of training epochs and hyperparameters. The mean and standard deviation are then computed from ten independent trials across three different random seeds. We subsequently validate our approach with real-time experiments on a physical robot. We designed three core tasks to evaluate the model’s ability to plan and execute diverse contact-rich behaviors, including exploit contact for stability and avoid contact for safety.
Task
(1) Support the Wall: the robot must resist external disturbances by stabilizing itself only through supportive hand contact; (2) Block the Ball: the robot must intercept a flying object only with defensive hand contact; (3) Traverse the Arch: the robot must pass through a low-clearance arch while avoiding unintended head contact.
Baselines
We compare our method with these baseline methods: (1) PPO: implementation in [2] . (2) ARWM: replace our framework with auto-regressive prediction training like [29, 30]. (3) Rew-MPC: replace objective function Eq. (16) with $\sum_{k=0}^{N-1}\gamma^{k}\hat{r}_{k}$ from PlaNet [24]. (4) TD-MPC: replace objective function Eq. (16) with $\sum_{k=0}^{N-1}\gamma^{k}\hat{r}_{k}+\gamma^{N}\hat{Q}$ from TD-MPC [23].
TABLE I: Single-Task Reward Evaluation of Our Method and Baselines: we analyze the influence of three key design choices on performance: the planning horizon NN, the world model training methodology, and the objective function.
METHOD Reward:Wall $\uparrow$ Reward:Ball $\uparrow$ Reward:Arch $\uparrow$
HORIZON NN
Ours, N=1 $0557\pm 0.0047 0.0$$066\pm 0.0050 0.0$$0396\pm 0.0121$
Ours, N=2 $0607\pm 0.0023 0.0$$056\pm 0.0003 0.0$$0154\pm 0.0011$
Ours, N=3 $0611\pm 0.0025 0.0$$059\pm 0.0003 0.0$$0156\pm 0.0011$
[…]
Ours, N=5 $0598\pm 0.0049 0.0$$058\pm 0.0012 0.0$$0144\pm 0.0062$
Ours, N=6 $\mathbf{0.0617}\pm 0.0031$ $0053\pm 0.0020 0.0$$0115\pm 0.0099$
WORLD MODEL
ARWM $0609\pm 0.0047 0.0$$039\pm 0.0033 0.0$$0018\pm 0.0183$
OBJECTIVE FUNCTION
Rew-MPC $2\pm 0.0204 0.0033±0.00$$3\pm 0.0044$ $0.0211\pm 0.0092$
TD-MPC $\mathbf{0.0699}\pm 0.0035$ $0016\pm 0.0047 0.0$$0145\pm 0.0005$
IV-A Advantages of the Use of Offline data
Sample Efficiency In Single-Task: We first compare our method against PPO implemented in [2], an online on-policy RL algorithm that remains the dominant training method in the legged robotics domain. A key distinction is that PPO requires continuous interaction with the environment, whereas our approach is fully offline, trained from a fixed, demonstration-free dataset without any environment interaction. We do not compare against off-policy methods such as SAC and its variants, which, although more sample-efficient than on-policy approaches, are less commonly applied to humanoid robots in real-world settings due to their limited scalability to complex, high-DoF dynamic control.
To compare single-task training performance, both our method and the PPO baseline leverage an identical low-level controller that tracks commands for end-effector position ($p_{ee}^{\top}$) and body height ($h_{body}$).
As shown in Fig. 5, in three contact-reward-dominant tasks, our method completes the task using only 0.5M data steps. In contrast, PPO requires a significantly larger amount of data, especially in simulations that necessitate visual rendering, where it consumes considerably more time. While PPO can quickly match our efficiency in tasks with simple visual features and a stationary robot, such as “Block the Ball,” our method achieves better performance on tasks with more complex visual representations and significant changes in the robot’s viewpoint. For instance, in the “Traverse the Arch” task, where the robot’s perspective changes dramatically between standing and squatting positions, our method substantially outperforms PPO.
Multi-Task Capability: In addition to sampling efficiency on the single task, we qualitatively analyze the challenges of applying online RL like PPO to a multi-task setting:
  • Reward Engineering: online RL requires either (1) a unified reward function across all tasks, which is difficult to design, or (2) substantial engineering effort to implement complex logic for task switching and conditional rewards.
  • Catastrophic Forgetting & Curriculum Design: online RL is prone to catastrophic forgetting. Mitigating this requires a carefully designed task sampling curriculum, the complexity of which grows as more tasks are added.
Our approach, which leverages offline data, circumvents these challenges by learning directly from a mixed dataset, enabling effective and scalable multi-task training. The result of our method in the multi-task setting is shown in Sec. IV-C.
Figure 7: Visualization of the world model’s prediction and planning process during the “Block the Ball” task. (a) The evolution of the Q-value map for the hand’s position in the X-Z plane. As the task unfolds (from Step 30 to 80, with the ball appearing at Step 60), the model dynamically updates its value estimates, with the high-value region (yellow) indicating the optimal location to intercept the object. (b) An open-loop prediction of future observations. Given an initial image, the model first reconstructs it (Horizon 0) and then generates a sequence of future frames (Horizon 2-16) by decoding its predicted latent states, visualizing its anticipation of the ball’s trajectory. It is worth noting that while our planner uses a shorter 4-step horizon, this visualization is for demonstrating the long-horizon physical coherence of our learned model.
IV-B Analysis of Key Design Choices
To demonstrate the necessity of our design, as shown in the Table I, we found that greedily maximizing the value (i.e., the horizon-1 case) is infeasible. This approach causes the robot to be myopic, favoring the maintenance of its default position and ignoring future contacts. We observe that longer horizons (e.g., N=6N{=}6) degrade performance, likely because bias dominates from longer-term prediction (see Sec. III-C), whereas N=4N{=}4 strikes a bias–variance sweet spot.
We also find that incorporating autoregressive prediction in our method (the ARWM baseline in Table I) is not necessary and can even be detrimental to value estimation in offline RL, as it overemphasizes precise prediction, which leads to value function overfitting.
As shown in the Table I, Rew-MPC, which uses reward as the objective function, yielded suboptimal results due to partial observability, noise, action lag, and sparse contact, which make rewards difficult to predict. And TD-MPC, which uses TD-target to evaluate trajectories, also costs unstable results. We attribute this to TD-error methods converging to deceptive solutions, where low TD error might mask highly inaccurate value estimates. As argued in [38], this phenomenon is caused by bias cancellation and the existence of infinitely many suboptimal solutions that satisfy the Bellman equation on an incomplete offline dataset.
IV-C Multi-Task Planning with Unified World Model
To evaluate the multi-task capability of our model, we trained a single model on a combined dataset from all tasks and compared it to the specialized single-task models from Table I. As shown in Fig. 6, the multi-task model achieves improved performance on two of the three tasks, with a minor drop in the “Block the Ball” task, which we attribute to its smaller reward scale.
To understand how this is achieved, we visualized the latent spaces using t-SNE. The visualizations reveal that our model learns to form distinct clusters for each task. The latent state $h_{t}$ shows significant temporal evolution, confirming that the model learns to encode the unique latent dynamics for each task from the mixed data.
IV-D Model Interpretation and Visualization on Prediction
We provide visualizations in Fig.7 that offer insight into the internal decision-making process of our framework. Specifically, we analyze this process on two levels: first, whether our model has learned a genuine understanding of the environment’s dynamics, and second, how it leverages this understanding for its decision-making process.
Long-Horizon Prediction: To verify the model’s ability to make long-horizon predictions, we visualize its open-loop rollouts. As illustrated in Fig. 7(b), given a single initial observation, the model first reconstructs it (Horizon 0) and then generates a sequence of expected future observations (Horizon 2-16) by decoding its predicted latent states. This predicted sequence clearly visualizes the model’s anticipation of the ball’s trajectory from left to right. This demonstrates that the model has learned a coherent and physically plausible model of the world, even without auto-regressive training and constrained solely by the joint-embedding loss. The blurriness in the generated frames is reasonable, as our highly compressed 32-dimensional latent space forces the model to extract only the most salient physical information
Contact-Directed Planning: Figure 7(a) visualizes how the model leverages its predictions for planning by showing the evolution of the objective function value in Eq. (11) for the hand’s target position. Early in the task (e.g., Step 30), the value map consistently encourages the hand to stay near a natural, energy-efficient default position. However, as the ball approaches and the plan solidifies (e.g., Step 60-70), a high-value region (yellow) emerges and sharpens, decisively guiding the robot’s hand toward the optimal contact point. This dynamic evolution of the value map showcases the model’s contact-directed reasoning, effectively forming an interpretable plan to achieve its objective.
IV-E Real-World Validation
We deployed our method for real-time experiments on Unitree G1 with 25 Hz real-time planning and evaluated a batch with 1024 action trajectories with a planning horizon of 4 steps at each timestep. The desired base velocity $v$ was controlled by a human operator.
Our real-world deployments, shown in Fig. 4, included both single-task and multi-task models, both of which proved capable of completing their assigned tasks. The policy also demonstrated the ability to generalize to out-of-distribution (OOD) scenarios, such as blocking a previously unseen box. These experiments validate that our method can achieve agile and robust vision-based control. Crucially, the learned policy exhibits reactive, context-dependent behavior rather than overfitting to a single action pattern. For instance, in the ’Support the Wall’ task, the robot only braces its hands against the wall when actively disturbed and returns to a neutral stance once balance is recovered. Please check our supplementary video for more details.
V Conclusion
By integrating a scalable ego-centric visual world model with value-guided sampling-based model predictive control, we demonstrate that humanoid robots can efficiently and robustly learn agile, contact-rich behaviors from offline, demonstration-free data, advancing data-efficient, vision-based planning for real-world robotic interaction.
ACKNOWLEDGMENT
This work was supported in part by NSF Grant CMMI-1944722 and by the Robotics and AI Institute. M. Ghaffari receives support by AFOSR MURI FA9550-23-1-0400. The authors would like to thank Jiaze Cai, Yen-Jen Wang for their help in experiments. We are also grateful to Bike Zhang, Fangchen Liu, Chaoyi Pan, Junfeng Long and Yiyang Shao for their valuable discussions. This work was done during H. Liu’s visit to UC Berkeley.